What is the directrix of the parabola 15y+3x^2=0
Answer by lwsshak3(11628)

What is the directrix of the parabola
15y+3x^2=0
3x^2=-15y
x^2=-5y
This is a parabola that opens down with vertex at the origin.
Its basic equation: x^2=-4py
For given equation:
4p=5
p=5/4
directrix: y=5/4